# On the Field Theoretical Description of an Alternative Model to Generalized Chaplygin Gas and its Thermodynamic Behaviour

### v1

@@ 4 Thermodynamic Analysis
Initial groundbreaking research by Songaila et al. [62] determined the cosmic microwave background temperature through neutral carbon atom measurements. Specifically, they estimated $T_{CMBR}=7.4\pm 0.8K$ at a redshift of $z=1.776$ in a cloud toward the quasar $Q1331+170$. Subsequent refined measurements by Mather et al. [63] established the present CMB black-body temperature at $T_{CMBR}=2.725\pm 0.002K$, measured locally at redshift $z=0$. Lima et al. [64] introduced an innovative perspective challenging standard cosmological models, proposing that the CMB temperature at high redshifts might deviate from predicted standard values. Their research incorporated novel concepts including decaying vacuum energy density and gravitational ’adiabatic’ photon creation, complemented by late inflationary models driven by scalar field dynamics. Srianand et al. [65] conducted a comprehensive investigation using high-resolution spectroscopy, imposing stringent upper limits on $T_{CMB}$ across a large sample of C I fine structure absorption lines. Their most notable finding came from detecting carbon monoxide in a damped Lyman-$\alpha$ system at $z_{abs}=2.41837$ toward SDSS $J143912.04+111740.5$, from which they determined $T_{CMBR}=9.15\pm 0.72K$. Building upon previous research, J. Ge et al. [66] identified absorption lines from ground and excited states of C I in the $z=1.9731$ damped Lyman-$\alpha$ system of QSO 0013-004. Through careful analysis of population ratios between excited and ground states, they estimated the CMBR temperature as $T=7.9\pm 1.0K$ at $0.61$ mm and $z=1.9731$, which aligned closely with contemporary standard cosmological predictions. Noterdaeme et al. [67] analysed a sub-damped Lyman-$\alpha$ system with neutral hydrogen column density at $z_{abs}=2.69$ toward SDSS $J123714.60+064759.5$. Their investigation revealed that CO excitation was predominantly influenced by radiative interaction with the CMBR, deriving $T_{ex}(CO)=10.5K$, corresponding to the expected $T_{CMBR}(z=2.69)=10.05K$. A systematic survey utilizing VLT/UVES spectrographs allowed P. Noterdaeme et al. [68] to constrain the CMBR temperature evolution to $z\approx 3$. By combining their measurements with previous constraints, they formulated a refined temperature model: $T_{CMB}(z)=(2.725\pm 0.002)\times(1+z)1-\beta K$, with $\beta=-0.007\pm 0.027$. These comprehensive investigations have collectively motivated further research into deriving precise expressions for the temperature of the Friedmann-Robertson-Walker (FRW) universe dominated by the proposed fluid matter as a function of redshift, aiming to rigorously validate existing cosmological models through meticulous observational constraints. We have presented our observational datasets in Tab. 1. The plot of the eq. (73) against the observational temperature values from CMB is presented in fig. 13. The plot shows the model is well fitted with the observed temperature values and the best fit model parameters are $T_{0}=2.725$, $V_{0}=0.845$, $\mu=0.889$ with the $\chi^{2}$/dof being $0.7$.
@@ 5 Validity of Generalised Second Law On Apparent Horizon and Cosmic Event Horizon
Now, we have deduced the eq. (98) using the eqs. (96), (97), (48) and (49). As, the expressions of $\dfrac{dS_{T}}{dt}$ and $T_{AH}$ is very complicated in form, so we do not present those in this paper, rather we have plotted them with respect to redshift $z$ in the following figures 14(b) and 14(c).
@@ 6.2 Results and Discussions
The plot 18(b) illustrates the redshift evolution of the dark energy equation of state parameter, $\omega(z)$, for the alternative model to the Generalized Chaplygin Gas (GCG) under different observational datasets: Pantheon+SH0ES (solid blue line), Union2.1 (dashed green line), and CC+BAO (dotted red line). A notable feature of the plot is the convergence of $\omega(z)$ values towards $-1$ as the redshift $z$ increases, indicating a transition towards a cosmological constant-like behaviour at high redshifts, which is consistent with the late-time accelerated expansion of the universe. At lower redshifts ($z\lesssim 1$), the three datasets exhibit distinguishable deviations, reflecting the sensitivity of the model parameters to different data constraints. Specifically, the Pantheon+SH0ES dataset suggests a slightly higher $\omega(z)$ at low $z$, whereas the CC+BAO dataset predicts a more negative $\omega(z)$, implying a stronger deviation from the cosmological constant scenario. This variation highlights the adaptability of the alternative model in capturing the observational nuances of dark energy dynamics. The model’s smooth evolution across all datasets further suggests its capability to provide a unified framework for understanding the interplay between dark energy and cosmic acceleration, particularly when tested against diverse observational constraints. These results underscore the robustness of the alternative model to GCG in addressing key cosmological challenges, including the Hubble tension, by effectively accommodating the observed differences in $\omega(z)$ while maintaining consistency with a broad range of data.
@@
The plot 18(a) depicts the evolution of the normalized energy density parameters, $\Omega(z)$, as functions of redshift $z$ for the alternative model. The green line represents the matter density parameter, $\Omega_{m}(z)$, while the blue line corresponds to the dark energy density parameter, $\Omega_{de}(z)$. The red dashed line marks the critical density, normalized to unity. At low redshifts ($z\sim 0$), the dark energy density dominates ($\Omega_{de}(z)>\Omega_{m}(z)$), consistent with the late-time acceleration of the universe. As $z$ increases, $\Omega_{m}(z)$ grows, reflecting the dominance of matter during earlier cosmic epochs, while $\Omega_{de}(z)$ diminishes, indicating the decreasing influence of dark energy in the high-redshift regime. Notably, the curves for $\Omega_{m}(z)$ and $\Omega_{de}(z)$ intersect at a transition redshift where the contributions of matter and dark energy are comparable, aligning with the expected behavior in standard cosmological models. The smooth behavior of both density parameters across redshift validates the robustness of the alternative model in describing the dynamic interplay between matter and dark energy components. These results highlight the model’s consistency with the standard cosmological paradigm while offering a framework to further probe deviations from $\Lambda$CDM at intermediate redshifts.
@@
    The DEC stipulates that the energy density dominates over the pressure magnitude, ensuring causal propagation of matter and energy. A violation of this condition can signal superluminal behavior, which is physically problematic in standard cosmology.
@@
The plots of SEC, DEC, and NEC against the redshift $z$ in Fig 19 reveal distinct behaviors for varying $\mu$. For $\mu=0.46$, the SEC exhibits a transition from positive to negative values, indicating a shift from decelerated to accelerated expansion. Similar trends are observed for $\mu=0.66$ and $\mu=0.86$, with all the transitions occurring at an redshift value $z\approx 0.5$. This suggests that the redshift value at which the violation of SEC occurs with the evolution of the universe, does not depend on the value of $\mu$ The NEC remains satisfied across the redshift range for all values of $\mu$, confirming the physical consistency of the model. In contrast, the DEC shows subtle deviations, particularly at lower redshifts for larger $\mu$ values. These deviations might imply that pressure contributions become comparable to or dominate the energy density in certain regimes. Overall, these results demonstrate that the alternative model to the GCG is consistent with the accelerated expansion of the universe while maintaining fundamental energy conditions. The parameter $\mu$ does not play an effective role in determining the transition epochs and the behavior of the energy conditions, offering a robust framework for understanding the interplay between dark energy and cosmic evolution.
@@
The alternative model to the generalized Chaplygin gas exhibits a set of key parameters that offer valuable insights into the cosmological dynamics of the universe. These parameters, evaluated against different datasets (Pantheon+SH0ES, Union2.1, and CC+BAO), are crucial in understanding the behavior of dark energy, matter, and the expansion of the universe. The matter density parameter at present time, $\Omega_{m0}$, varies across different datasets. For the Pantheon+SH0ES dataset, the best-fit value is $\Omega_{m0}=0.237^{+0.081}_{-0.057}$, while for Union2.1 and CC+BAO, the values are $\Omega_{m0}=0.252^{+0.066}_{-0.037}$ and $\Omega_{m0}=0.290^{+0.077}_{-0.055}$, respectively. These variations indicate slight shifts in the matter density as constrained by different observational datasets, suggesting that the model is sensitive to the choice of data but still remains within the range expected from current cosmological observations [101, 89, 98, 100]. The equation of state parameter for matter, $\omega_{m}$, describes the behaviour of matter under the alternative model. For Pantheon+SH0ES, $\omega_{m}=-0.033^{+0.091}_{-0.047}$, for Union2.1, $\omega_{m}=0.057^{+0.170}_{-0.116}$, and for CC+BAO, $\omega_{m}=-0.067^{+0.039}_{-0.022}$. These values suggest that the matter component is relatively close to the standard value of $\omega_{m}=0$ (indicating pressureless dust), though slight deviations are present, which reflect the model’s flexibility in accommodating different scenarios with the presence of some exotic type matter. The parameter $\mu$, which plays a role in modifying the equation of state, is found to be consistent across all datasets. The values are $\mu=0.863^{+0.048}_{-0.067}$ for Pantheon+SH0ES, $\mu=0.865^{+0.062}_{-0.086}$ for Union2.1, and $\mu=0.841^{+0.076}_{-0.088}$ for CC+BAO. The consistency of this parameter indicates a relatively stable behavior of dark energy under the alternative model, independent of the dataset used and consistent with the chosen value from ref [47] and with the observationally fitted values [48, 49] which is presented in Sect. 2. The parameter $V_{0}$, which could be associated with the normalization of the volume term, shows a noticeable variation across the datasets. For Pantheon+SH0ES, $V_{0}=1.028^{+0.234}_{-0.278}$, for Union2.1, $V_{0}=0.639^{+0.294}_{-0.331}$, and for CC+BAO, $V_{0}=0.607^{+0.189}_{-0.302}$. These differences reflect the sensitivity of the model to the observational constraints and the underlying cosmological parameters. The age of the universe is also an important parameter that has been constrained by the alternative model. The best-fit age is $14.04^{+0.14}_{-0.12}$ Gyr, which aligns well with the standard cosmological model [101] with a slight larger value, further validating the model’s consistency with the observed age of the universe.

### v2

@@ 4 Thermodynamic Analysis
Initial groundbreaking research by Songaila et al. [63] determined the cosmic microwave background temperature through neutral carbon atom measurements. Specifically, they estimated $T_{CMBR}=7.4\pm 0.8K$ at a redshift of $z=1.776$ in a cloud toward the quasar $Q1331+170$. Subsequent refined measurements by Mather et al. [64] established the present CMB black-body temperature at $T_{CMBR}=2.725\pm 0.002K$, measured locally at redshift $z=0$. Lima et al. [65] introduced an innovative perspective challenging standard cosmological models, proposing that the CMB temperature at high redshifts might deviate from predicted standard values. Their research incorporated novel concepts including decaying vacuum energy density and gravitational ’adiabatic’ photon creation, complemented by late inflationary models driven by scalar field dynamics. Srianand et al. [66] conducted a comprehensive investigation using high-resolution spectroscopy, imposing stringent upper limits on $T_{CMB}$ across a large sample of C I fine structure absorption lines. Their most notable finding came from detecting carbon monoxide in a damped Lyman-$\alpha$ system at $z_{abs}=2.41837$ toward SDSS $J143912.04+111740.5$, from which they determined $T_{CMBR}=9.15\pm 0.72K$. Building upon previous research, J. Ge et al. [67] identified absorption lines from ground and excited states of C I in the $z=1.9731$ damped Lyman-$\alpha$ system of QSO 0013-004. Through careful analysis of population ratios between excited and ground states, they estimated the CMBR temperature as $T=7.9\pm 1.0K$ at $0.61$ mm and $z=1.9731$, which aligned closely with contemporary standard cosmological predictions. Noterdaeme et al. [68] analysed a sub-damped Lyman-$\alpha$ system with neutral hydrogen column density at $z_{abs}=2.69$ toward SDSS $J123714.60+064759.5$. Their investigation revealed that CO excitation was predominantly influenced by radiative interaction with the CMBR, deriving $T_{ex}(CO)=10.5K$, corresponding to the expected $T_{CMBR}(z=2.69)=10.05K$. A systematic survey utilizing VLT/UVES spectrographs allowed P. Noterdaeme et al. [69] to constrain the CMBR temperature evolution to $z\approx 3$. By combining their measurements with previous constraints, they formulated a refined temperature model: $T_{CMB}(z)=(2.725\pm 0.002)\times(1+z)1-\beta K$, with $\beta=-0.007\pm 0.027$. These comprehensive investigations have collectively motivated further research into deriving precise expressions for the temperature of the Friedmann-Robertson-Walker (FRW) universe dominated by the proposed fluid matter as a function of redshift, aiming to rigorously validate existing cosmological models through meticulous observational constraints. We have presented our observational datasets in Tab. 2. The plot of the eq. (73) against the observational temperature values from CMB is presented in fig. 15. The plot shows the model is well fitted with the observed temperature values and the best fit model parameters are $T_{0}=2.725$, $V_{0}=0.845$, $\mu=0.889$ with the $\chi^{2}$/dof being $0.7$. The comparison with CMB data serves as a crucial test for the viability of the proposed model. The key insights derived from this comparison include the consistency with the observed thermal history, where the predicted temperature evolution aligns with CMB temperature measurements, validating the model’s compatibility with early universe thermodynamics. This agreement indicates that modified gravity effects do not disrupt the standard thermal history during the recombination era. Additionally, while the primary CMB spectrum remains consistent with standard predictions, subtle deviations in higher-order acoustic peaks suggest modifications to the cosmic fluid’s sound speed and thermal diffusion properties, providing indirect evidence of the fluid’s altered thermodynamic properties. Moreover, the observational comparison imposes constraints on key model parameters, such as the coupling strength of scalar fields and the modification scale of gravitational interactions, ensuring that the proposed phase transitions occur at cosmologically relevant epochs without conflicting with CMB measurements. Methodologically, the study advances the field by integrating thermodynamic analysis with observational constraints from CMB data, bridging the gap between theoretical predictions and empirical observations. This combined approach identifies thermodynamic signatures as observational probes, enabling future CMB experiments and large-scale structure surveys to test the proposed phase transitions and stability conditions. Furthermore, by incorporating thermodynamic constraints, the study refines the parameter space of viable dark energy and modified gravity models, enhancing their predictive power. The approach is also extendable to other cosmological datasets, such as baryon acoustic oscillations (BAO) and supernovae data, providing a robust framework for testing the thermodynamic consistency of alternative cosmological models.
@@
A key aspect of our analysis involves checking the stability of the proposed alternative fluid model under classical perturbations. This is done by computing the squared speed of sound, $v_{s}^{2}=\frac{dp}{d\rho}$. In our case, we find that $v_{s}^{2}$ remains positive throughout cosmic evolution (see Fig. 11(a)), implying that the model is free from ghost instabilities. Additionally, the specific heat at constant volume, $C_{V}$, is computed to verify the thermodynamic stability of the model. Our findings indicate that the entropy of the system is always increasing, thus this must satisfy the Generalized Second Law (GSL) at the apparent horizon, though violations can occur at the event horizon at certain epochs. This suggests that the choice of horizon significantly affects the thermodynamic consistency of the model, a feature not commonly explored in previous works on Chaplygin-like fluid models.
@@ 5 Validity of Generalised Second Law On Apparent Horizon and Cosmic Event Horizon
Now, we have deduced the eq. (98) using the eqs. (96), (97), (48) and (49). As, the expressions of $\dfrac{dS_{T}}{dt}$ and $T_{AH}$ is very complicated in form, so we do not present those in this paper, rather we have plotted them with respect to redshift $z$ in the following figures 16(b) and 16(c).
@@
The study is conducted for two horizons: the cosmological event horizon and the apparent horizon. For the event horizon, the entropy rate is determined using Eq. (91) and plotted in Fig. 16(a), indicating GSL violation during certain evolutionary epochs but ensuring future satisfaction. For the apparent horizon, using Eq. (98), Fig. 16(b) demonstrates that GSL is always satisfied. Additionally, the Kodama-Hayward temperature plotted in Fig. 16(c) remains positive throughout, ensuring thermodynamic stability.
The sign of $\dfrac{dS_{T}}{dT}$ plays a crucial role in determining the stability and physical nature of cosmic evolution. At the apparent horizon, where $\dfrac{dS_{T}}{dT}$ is positive, entropy increases with time, indicating that the system tends to evolve toward a more thermodynamically stable state. This positive gradient suggests that perturbations or deviations from equilibrium are naturally damped, contributing to the long-term stability of the cosmological model. Conversely, at the cosmological event horizon, where $\dfrac{dS_{T}}{dT}$ is negative, the entropy decreases with time. This implies that the system may evolve away from equilibrium, making it more susceptible to instabilities or non-equilibrium effects. Physically, this reflects the influence of the accelerating expansion of the universe, where the event horizon acts as a boundary that limits causal interactions, potentially affecting the entropy flow and thermodynamic balance. It is important to emphasize that the analysis of GSL not only examines the thermodynamic evolution of the universe but also provides insight into the stability of the underlying cosmological model. The positivity of both $\dfrac{dS_{T}}{dt}$ and $T_{AH}$ indicates that the model is thermodynamically stable at the apparent horizon, as any deviation would imply instability. In contrast, the negative entropy gradient at the cosmological event horizon highlights the complex interplay between accelerated expansion and thermodynamic evolution, suggesting that additional factors may influence long-term stability. Therefore, explicit considerations of stability are crucial, as thermodynamic stability ensures the robustness and physical viability of the background theory. Hence, the consistency observed in these results strengthens the validity of the proposed cosmological model, aligning with the fundamental principles of thermodynamics and cosmic evolution.
@@ 6.2 Results and Discussions
The plot 20(b) illustrates the redshift evolution of the dark energy equation of state parameter, $\omega(z)$, for the alternative model to the Generalized Chaplygin Gas (GCG) under different observational datasets: Pantheon+SH0ES (solid blue line), Union2.1 (dashed green line), and CC+BAO (dotted red line). A notable feature of the plot is the convergence of $\omega(z)$ values towards $-1$ as the redshift $z$ increases, indicating a transition towards a cosmological constant-like behaviour at high redshifts, which is consistent with the late-time accelerated expansion of the universe. At lower redshifts ($z\lesssim 1$), the three datasets exhibit distinguishable deviations, reflecting the sensitivity of the model parameters to different data constraints. Specifically, the Pantheon+SH0ES dataset suggests a slightly higher $\omega(z)$ at low $z$, whereas the CC+BAO dataset predicts a more negative $\omega(z)$, implying a stronger deviation from the cosmological constant scenario. This variation highlights the adaptability of the alternative model in capturing the observational nuances of dark energy dynamics. The model’s smooth evolution across all datasets further suggests its capability to provide a unified framework for understanding the interplay between dark energy and cosmic acceleration, particularly when tested against diverse observational constraints. These results underscore the robustness of the alternative model to GCG in addressing key cosmological challenges, including the Hubble tension, by effectively accommodating the observed differences in $\omega(z)$ while maintaining consistency with a broad range of data.
@@
The plot 20(a) depicts the evolution of the normalized energy density parameters, $\Omega(z)$, as functions of redshift $z$ for the alternative model. The green line represents the matter density parameter, $\Omega_{m}(z)$, while the blue line corresponds to the dark energy density parameter, $\Omega_{de}(z)$. The red dashed line marks the critical density, normalized to unity. At low redshifts ($z\sim 0$), the dark energy density dominates ($\Omega_{de}(z)>\Omega_{m}(z)$), consistent with the late-time acceleration of the universe. As $z$ increases, $\Omega_{m}(z)$ grows, reflecting the dominance of matter during earlier cosmic epochs, while $\Omega_{de}(z)$ diminishes, indicating the decreasing influence of dark energy in the high-redshift regime. Notably, the curves for $\Omega_{m}(z)$ and $\Omega_{de}(z)$ intersect at a transition redshift where the contributions of matter and dark energy are comparable, aligning with the expected behaviour in standard cosmological models. The smooth behaviour of both density parameters across redshift validates the robustness of the alternative model in describing the dynamic interplay between matter and dark energy components. These results highlight the model’s consistency with the standard cosmological paradigm while offering a framework to further probe deviations from $\Lambda$CDM at intermediate redshifts.
@@
    The DEC stipulates that the energy density dominates over the pressure magnitude, ensuring causal propagation of matter and energy. A violation of this condition can signal superluminal behaviour, which is physically problematic in standard cosmology.
@@
The plots of SEC, DEC, and NEC against the redshift $z$ in Fig 21 reveal distinct behaviours for varying $\mu$. For $\mu=0.46$, the SEC exhibits a transition from positive to negative values, indicating a shift from decelerated to accelerated expansion. Similar trends are observed for $\mu=0.66$ and $\mu=0.86$, with all the transitions occurring at an redshift value $z\approx 0.5$. This suggests that the redshift value at which the violation of SEC occurs with the evolution of the universe, does not depend on the value of $\mu$ The NEC remains satisfied across the redshift range for all values of $\mu$, confirming the physical consistency of the model. In contrast, the DEC shows subtle deviations, particularly at lower redshifts for larger $\mu$ values. These deviations might imply that pressure contributions become comparable to or dominate the energy density in certain regimes. Overall, these results demonstrate that the alternative model to the GCG is consistent with the accelerated expansion of the universe while maintaining fundamental energy conditions. The parameter $\mu$ does not play an effective role in determining the transition epochs and the behaviour of the energy conditions, offering a robust framework for understanding the interplay between dark energy and cosmic evolution.
@@
The alternative model to the generalized Chaplygin gas exhibits a set of key parameters that offer valuable insights into the cosmological dynamics of the universe. These parameters, evaluated against different datasets (Pantheon+SH0ES, Union2.1, and CC+BAO), are crucial in understanding the behaviour of dark energy, matter, and the expansion of the universe. The matter density parameter at present time, $\Omega_{m0}$, varies across different datasets. For the Pantheon+SH0ES dataset, the best-fit value is $\Omega_{m0}=0.237^{+0.081}_{-0.057}$, while for Union2.1 and CC+BAO, the values are $\Omega_{m0}=0.252^{+0.066}_{-0.037}$ and $\Omega_{m0}=0.290^{+0.077}_{-0.055}$, respectively. These variations indicate slight shifts in the matter density as constrained by different observational datasets, suggesting that the model is sensitive to the choice of data but still remains within the range expected from current cosmological observations [102, 90, 99, 101]. The equation of state parameter for matter, $\omega_{m}$, describes the behaviour of matter under the alternative model. For Pantheon+SH0ES, $\omega_{m}=-0.033^{+0.091}_{-0.047}$, for Union2.1, $\omega_{m}=0.057^{+0.170}_{-0.116}$, and for CC+BAO, $\omega_{m}=-0.067^{+0.039}_{-0.022}$. These values suggest that the matter component is relatively close to the standard value of $\omega_{m}=0$ (indicating pressureless dust), though slight deviations are present, which reflect the model’s flexibility in accommodating different scenarios with the presence of some exotic type matter. The parameter $\mu$, which plays a role in modifying the equation of state, is found to be consistent across all datasets. The values are $\mu=0.863^{+0.048}_{-0.067}$ for Pantheon+SH0ES, $\mu=0.865^{+0.062}_{-0.086}$ for Union2.1, and $\mu=0.841^{+0.076}_{-0.088}$ for CC+BAO. The consistency of this parameter indicates a relatively stable behaviour of dark energy under the alternative model, independent of the dataset used and consistent with the chosen value from ref [45] and with the observationally fitted values [53, 54] which is presented in Sect. 2. The parameter $V_{0}$, which could be associated with the normalization of the volume term, shows a noticeable variation across the datasets. For Pantheon+SH0ES, $V_{0}=1.028^{+0.234}_{-0.278}$, for Union2.1, $V_{0}=0.639^{+0.294}_{-0.331}$, and for CC+BAO, $V_{0}=0.607^{+0.189}_{-0.302}$. These differences reflect the sensitivity of the model to the observational constraints and the underlying cosmological parameters. The age of the universe is also an important parameter that has been constrained by the alternative model. The best-fit age is $14.04^{+0.14}_{-0.12}$ Gyr, which aligns well with the standard cosmological model [102] with a slight larger value, further validating the model’s consistency with the observed age of the universe.
The comparison of our alternative fluid model with the standard $\Lambda$CDM model using recent observational data, including DESI and DESY5, demonstrates that both models provide a good fit to the data. The best-fit parameters and information criteria, presented in Tables 3 and 4, indicate that while $\Lambda$CDM yields slightly lower AIC, BIC, and DIC values, the differences are not statistically significant. This suggests that our model remains competitive, particularly considering the potential for evolving dark energy hinted at by recent DESI and DESY5 analyses. The DESI data, with its precise measurements of baryon acoustic oscillations (BAO), and the DESY5 supernova compilation, which offers improved constraints on cosmic expansion, both align closely with the predictions of a cosmological constant. However, subtle deviations observed in some analyses suggest the possibility of time-varying dark energy [103, 104]. Our model, characterized by the parameters $\mu$ and $V_{0}$, accommodates such dynamics while maintaining consistency with current observational constraints. Although the differences in information criteria do not strongly favour our model, its ability to reproduce the observed thermal and expansion histories of the universe while allowing for potential deviations from $\Lambda$CDM makes it a viable alternative. In conclusion, the incorporation of DESI and DESY5 data reinforces the robustness of our fluid model as a realistic alternative to $\Lambda$CDM. The consistency of its predictions with observational constraints and the possibility of capturing evolving dark energy dynamics highlight its potential for further exploration as more precise cosmological data become available.